# Retrieval-Oriented Knowledge for Click-Through Rate Prediction

### v1

@@ 6. Industry Application
In this section, we will demonstrate the superior performance of ROK in training and inference efficiency.
@@ 6.1. Training Efficiency
In evaluating model efficiency, it’s vital to look at both inference speed and training time. Training ROK involves pre-training a sample-level retrieval-based model, creating retrieval-oriented knowledge, and applying this knowledge. Despite these steps, the extra training time is reasonable compared to a standalone model. The training duration is detailed in Table 5. Experiments were run on an AMD EPYC 7T83 with an RTX 4090, using hyperparameters—learning rate, batch size, weight decay—optimized via grid search as per the RIM study. We also implemented early stopping after 3 steps to enhance efficiency. Two observations stand out. Firstly, incorporating ROK into the backbone model significantly reduces training time, thanks to fewer parameters needing updates, partly because ROK’s parameters are frozen and the embedding size is halved. This leads to faster convergence. Second, the total time of Phase 1 and Phase 2 is close to the training duration of the DIEN without ROK. The relatively time-consuming part of the whole training process would be the RIM’s retrieval mechanism. However, this is mitigated by using large bulk sizes and caching, which decreases retrieval times. For further efficiency, extending the knowledge base update frequency beyond the backbone model’s could be beneficial.
@@ 6.2. Inference Efficiency
In the real world of online services, from advertising platforms to recommendation systems, the significance of inference efficiency cannot be ignored. The introduction of the knowledge base significantly amplifies the efficiency in sample-level retrieval, trimming down the time complexity from $\mathcal{O}(N\log\left(N\right))$, where $N$ is the search pool size to a constant time $\mathcal{O}(1)$ through bypassing conventional retrieval & aggregation mechanisms.
Within this context, we thoroughly assess the inference efficiency of various backbone models—DeepFM (Guo et al., 2017), DIEN (Zhou et al., 2019), and DIN (Zhou et al., 2018)—when seamlessly integrated with ROK and juxtaposed against retrieval-based methods such as UBR (Qin et al., 2020), RIM (Qin et al., 2021), and DERT (Zheng et al., 2023). The comparisons drawn from our analyses, as detailed in Table 6, offer insights. Specifically, due to the $\mathcal{O}(N\log\left(N\right))$ complexity of the online retrieval process, the inference time of retrieval-based methods is significantly longer than other baseline models. For instance, RIM’s inference time is an overwhelming $1\sim 2$ orders of magnitude longer than DeepFM, DIEN, and DIN. This significant delay, despite any performance enhancements, fails to align with the real-time response expectations of RSs. These systems necessitate a swift response, making such latencies of retrieval-based methods industrially infeasible. Moreover, deploying retrieval-based methods for real-world recommendation systems remains complex and costly due to challenges such as encoding discretized features to dense representations, high resource consumption, and divergent inference processes. In contrast, our innovative ROK distinguishes itself by delivering notable improvements over the backbone models while maintaining negligible additional inference time, making sample-level retrieval-based methods feasible in online service deployment. As shown in Table 6, ROK consistently improves the AUC of backbone models across all datasets while only slightly increasing the inference speed by 0.12ms to 0.25ms per sample. Compared to retrieval-based methods, ROK achieves comparable or even better performance with significantly faster inference speed. For example, on the Tmall dataset, DIEN+ROK surpasses the performance of DERT (0.9226 vs. 0.9200) while being more than 3 times faster in inference (5.51ms vs. 17.78ms per sample).
ROK’s design allows it to be retrieval-free and compatible with existing recommendation systems. It only needs to perform inference once, avoiding additional data transfer overhead and making it highly efficient. This represents a blend of predictive accuracy and rapid inference, making ROK a noteworthy solution for modern online services.
Table 6. Comparison of AUC and inference speed (ms per sample) across models on Tmall, Taobao, and Alipay.
@@
AUC Inference Speed AUC Inference Speed AUC Inference Speed
DeepFM 0.8585 1.34 0.6710 1.36 0.6971 1.19
DeepFM+ROK 0.8946 1.46 0.6946 1.40 0.7361 1.41
DIEN 0.8839 5.44 0.7506 4.89 0.7485 3.69
DIEN+ROK 0.9226 5.51 0.8382 5.04 0.7884 3.73
DIN 0.8796 1.28 0.7433 1.32 0.7647 1.43
DIN+ROK 0.9171 1.46 0.8091 1.36 0.8093 1.56
UBR (Retrieval) 0.8975 20.71 (17.67) 0.8169 56.45 (53.32) 0.7952 30.32 (27.31)
RIM (Retrieval) 0.9151 174.81 (173.27) 0.8567 206.22 (204.78) 0.8005 113.95 (112.37)
DERT (Retrieval) 0.9200 17.78 (16.19) 0.8647 19.53 (16.97) 0.8087 17.95 (16.31)
@@ 7. Conclusion
This paper introduces the ROK framework, addressing the inefficiency challenges inherent in sample-level retrieval-based methods during inference. By developing a neural network-based knowledge base, ROK imitates and preserves the essential retrieval representations through a novel decomposition-reconstruction approach, further optimized via knowledge distillation and contrastive learning. This allows for the integration of retrieval-oriented knowledge with various CTR models, either at an instance or feature level. Through rigorous testing on three public datasets, ROK not only matches the performance of existing retrieval-based CTR models but also demonstrates superior inference efficiency. These findings highlight ROK’s potential to significantly improve the efficiency and effectiveness of CTR models.
Appendix A Appendix
A.1. T-sne visualization
Figure 7 depict the knowledge distribution patterns of both ROK and RIM in Tmall dataset.
Refer to captionRefer to caption
((a)) Tmall (Train)
Refer to captionRefer to caption
((b)) Tmall (Test)
Figure 7. The t-SNE visualization of knowledge from RIM (on the left in each subfigure) and ROK (on the right in each subfigure) for Tmall.
A.2. Ablation & Hyper-parameter Study on Taobao
In this section, we present the ablation and hyper-parameter study conducted on the Taobao dataset. The results are depicted in Figure 8, Figure 9.
Figure 8. Comparison of AUC scores for different update strategies on Taobao.
Figure 9. Comparison of AUC scores across varying $\alpha$ values on Taobao. At $\alpha=0$, only Retrieval Imitation loss is utilized, whereas at $\alpha=1$, only the Contrastive Regularization loss is applied.

### v2

@@ 6.1. Training Efficiency
Training ROK involves pre-training a sample-level retrieval-based model, creating retrieval-oriented knowledge, and applying this knowledge to backbone CTR models. Despite these steps, the extra training time is reasonable compared to training a standalone model. The training duration is detailed in Table 5. Two observations stand out. Firstly, incorporating ROK into the backbone model significantly reduces training time, thanks to fewer parameters needing updates, partly because ROK’s parameters are frozen and the embedding size is halved. This leads to faster convergence. Second, the total time of Phase 1 and Phase 2 is close to the training duration of the DIEN without ROK. The relatively time-consuming part of the whole training process would be the RIM’s retrieval mechanism. However, this is mitigated by using large bulk sizes and caching, which decreases retrieval times. For further efficiency, extending the knowledge base update frequency beyond that of the backbone model could be beneficial.
@@ 6.2. Inference Efficiency
Deploying sample-level retrieval-based methods is complex and costly due to several key factors like extreme inefficiency problems during inference, high resource consumption and divergent inference process that requires an inference-retrieval-inference cycle. Our method, ROK, introduces a knowledge base that improves retrieval efficiency, reducing time complexity to $\mathcal{O}(1)$. We assessed backbone models—DeepFM, DIEN, and DIN—integrated with ROK and compared them to UBR (Qin et al., 2020), sparse sample-level retrieval-based methods, and dense sample-level retrieval-based methods. As shown in Table 6, retrieval-based methods have significantly longer inference times, making them impractical for real-time requirements. In contrast, ROK improves backbone models with minimal additional inference time, making it suitable for online deployment. Table 3 and Table 6 demonstrate that ROK consistently improves the AUC of all backbone models across the tested datasets while only slightly increasing the inference time by 0.12 ms to 0.25 ms per sample, showcasing its ability to enhance performance without significantly impacting speed.
ROK’s retrieval-free design ensures compatibility with existing systems, requiring only one inference process, thus avoiding data transfer overhead and enhancing efficiency. This combination of accuracy and rapid inference makes ROK a superior solution for modern online services.
Table 6. Comparison of inference speed (ms per sample) across models on Tmall, Taobao, and Alipay. For UBR, sparse retrieval and dense retrieval, both total inference and retrieval times are presented to highlight the impact of the retrieval process.
@@
UBR (Retrieval) 20.71 (17.67) 56.45 (53.32) 30.32 (27.31)
Sparse (Retrieval) 174.81 (173.27) 206.22 (204.78) 113.95 (112.37)
Dense (Retrieval) 17.78 (16.19) 19.53 (16.97) 17.95 (16.31)
DeepFM 1.34 1.36 1.19
DeepFM+ROK 1.46 1.40 1.41
DIEN 5.44 4.89 3.69
DIEN+ROK 5.51 5.04 3.73
DIN 1.28 1.32 1.43
DIN+ROK 1.46 1.36 1.56
@@ 7. Conclusion
The ROK framework tackles the critical challenges faced by sample-level retrieval-based methods, such as inference inefficiency and high resource consumption. By introducing a neural network-based Knowledge Base, ROK effectively distills and stores retrieval-oriented knowledge from pre-trained retrieval-based models. This innovative approach enables the seamless integration of retrieval-enhanced representations with various CTR models at both the instance and feature levels, bypassing the need for time-consuming retrieval during inference. Extensive experiments on three large-scale datasets demonstrate ROK’s exceptional compatibility and performance, significantly enhancing the performance of various CTR methods. Remarkably, the introduction of contrastive regularization allows ROK to surpass the performance of the sample-level retrieval-based teacher model, highlighting its strong potential for real-world applications. The ROK framework successfully transforms sample-level retrieval-based methods into practical, efficient, and scalable solutions for industrial deployment. By eliminating the need for resource-intensive retrieval during inference and demonstrating the feasibility of distilling knowledge from non-parametric models using a parametric approach, ROK opens up new possibilities for leveraging vast amounts of historical data in CTR prediction.
8. Acknowledgement
We thank MindSpore (min, 2020) for the partial support of this work, which is a new deep learning computing framework.